### v1

# CCFC: Bridging Federated Clustering and Contrastive Learning

Abstract
Federated clustering, an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data. However, the combination of federated clustering and representation learning remains underexplored. To bridge this, we first tailor a cluster-contrastive model for learning clustering-friendly representations. Then, we harness this model as the foundation for proposing a new federated clustering method, named cluster-contrastive federated clustering (CCFC). Benefiting from representation learning, the clustering performance of CCFC even double those of the best baseline methods in some cases. Compared to the most related baseline, the benefit results in substantial NMI score improvements of up to 0.4155 on the most conspicuous case. Moreover, CCFC also shows superior performance in handling device failures from a practical viewpoint.
1 Introduction
Federated clustering, an essential extension of centralized clustering for federated scenarios, enables multiple data-holding clients to collaboratively group data while keeping their data locally [8, 27, 31, 30]. It often constitutes a significant initial step in many learning tasks, such as client-selection [10] and personalization [20, 6]. And it is natural to extend centralized clustering methodologies into federated scenarios. In centralized scenarios, clustering driven by representation learning has made significant advancements in handling high-dimensional complex data, particularly images [34]. However, the marriage of federated clustering and representation learning (e.g. contrastive learning [24, 2]) remains an underexplored research avenue.
(a) Raw pixels
(b) Standalone local model
(c) FedAvg global model
(d) FedAvg local model
Figure 1: t-SNE visualizations on MNIST (best viewed in color). (a) shows the local data in the original data space, where each color corresponds to a specific category of digits in MNIST. (b) - (c) show the local data in different latent spaces.
To illustrate the necessity of the marriage and the challenges it poses, we simulate a simple federated scenario on MNIST, where images are evenly distributed among 10 clients. Fig. 0(a) showcases the t-SNE visualization of a randomly selected local dataset in its original data space, and one can observe that many samples from different categories are mixed up together, necessitating representation learning. To handle this, there are two straightforward methods: 1) Independent execution of a SimSiam model [5] by each client; 2) Integration of the SimSiam model into the Federated Averaging (FedAvg) framework [21]. As shown in Fig. 0(b) and Fig. 0(c), one can see that: 1) The local model trained solely on local data exhibits poor performance. 2) The collaborative training yields a superior global model. However, in the subsequent round of local updates, the limited information from the local data can lead to model regression (Fig. 0(d)). A similar problem of the model regression has also been observed in federated classification tasks, and it can be alleviated by introducing a model-contrastive loss to regularize the local model to refrain from deviating excessively from the global model during the local training [17].
Motivated by these, we first tailor a contrastive model, named cluster-contrastive model, for learning clustering-friendly representations. Then, we harness this model as the foundation for proposing a simple but effective federated clustering method, named cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term. Throughout the entire training procedure of CCFC, the only shared information between clients and the server is cluster-contrastive models and cluster centroids, thereby preserving data privacy. In each communication round, CCFC involves three main steps: global information dissemination, local training and local information aggregation. In the first step, each client first downloads a global model and $k$ global cluster centroids from the server, and then update the local model with the downloaded one, where $k$ is the number of clusters. In the second step, each client performs cluster assignment by labeling their local data with the index of the nearest global cluster centroid in the latent space, and they train the local model using both the local data and the assignment results. After this, k-means (KM) [19] is employed to create $k$ local cluster centroids for storing local semantic information. In the final step, clients upload their trained local models and local cluster centroids to the server, which aggregates these local models into a new global model through weighted averaging, and aggregates these local cluster centroids into $k$ new global cluster centroids through KM.
Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. The most related method to our CCFC is k-fed [8], with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI [28] score improvements of up to 0.4155 and Kappa [18] score enhancements of 0.4593 on the most conspicuous case. In summary, our contributions are as follows:
  • A tailored cluster-contrastive model is introduced to facilitate the acquisition of more clustering-friendly representations.
  • Leveraging this model, a simple yet effective federated clustering method, CCFC, is proposed.
  • Comprehensive experiments validate the significant superiority of CCFC and offer valuable insights into its performance.
2 Related Work
2.1 Clustering
Clustering, a fundamental machine learning task, aims to group similar samples together and often constitutes a significant initial step in many centralized or federated learning tasks, including information retrieval [33, 1], anomaly detection [14, 15], client-selection [10] and personalization [20, 6]. Traditionally, it is assumed that data is consolidated on a central server for model training. However, in federated scenarios, data is distributed among multiple isolated clients, and the sharing of local data is prohibited due to privacy concerns. In this scenario, clustering tasks are fraught with significant difficulties, as local data alone is inadequate for the accurate clustering of itself, and a centralized global dataset remains unfeasible [27].
To address these, federated clustering (FC) has arisen, allowing multiple clients to collaboratively group data while preserving their data locally [8, 27, 31, 30]. As an extension of centralized clustering, FC inherently embodies an exploratory path, i.e. extending centralized clustering methodologies to federated scenarios. Representative extensions include k-FED [8] and SDA-FC-KM [31], derived from k-means [19]; FFCM [27] and SDA-FC-FCM [31], extensions of fuzzy c-means [3]; and PPFC-GAN [30], an outgrowth of DCN [32]. Despite these extensions have advanced FC, their progress is limited. For instance, on the MNIST dataset, the NMI [28] values of the top 10 centralized clustering methods surpass 0.93111https://paperswithcode.com/sota/image-clustering-on-mnist-full. In stark contrast, even in the simplest federated scenarios, where local data among clients is independently-and-identically-distributed (IID), the NMI value of the cutting-edge PPFC-GAN falls short of 0.7 (Tab. 1).
In the realm of centralized clustering, the rapid progress can largely be attributed to the incorporation of representation learning techniques [34]. Hence, it stands to reason that federated clustering methods powered by representation learning may hold the key to narrowing this existing gap. However, the combination of federated clustering and representation learning remains underexplored.
2.2 Contrastive Learning
Contrastive learning [24, 2] has emerged as a prominent paradigm within the realm of representation learning, particularly for the acquisition of visual representations. While contrastive learning has been extensively explored and proven effective in centralized scenarios [9, 26, 22], its extension into federated scenarios has been relatively limited. These extensions can be broadly categorized into two groups: one employs contrastive learning to handle the challenges posed by the non-IID problem in federated classification tasks [17, 29, 23], while the other harnesses it for the acquisition of generic representations for downstream learning tasks [35, 12, 36].
As far as we can ascertain, no extensions have been introduced to tackle federated clustering tasks. To bridge contrastive learning and federated clustering, we tailor a new contrastive model for clustering tasks, called cluster-contrastive model. Leveraging this model, we propose a simple yet effective federated deep clustering method named cluster-contrastive federated clustering (CCFC).
Figure 2: Cluster-contrastive model architecture. $n_{c}$ images from cluster $c$ are encoded by the same encoder $f$, which comprises a backbone (e.g. ResNet-18) and an MLP. Then, an MLP predictor is applied to the $i$-th path, while $n_{c}-1$ stop-gradient operations are applied to the remaining paths. The model maximizes the average similarity of $x_{i,c}$ to other images.
3 Cluster-Contrastive Federated Clustering
3.1 Cluster-Contrastive Model
Given $k$ clusters obtained from clustering, each one comprises $n_{c}$ images. As shown in Fig. 2, the proposed model takes $n_{c}$ images from cluster $c$ as inputs. These images are processed by the same encoder $f$, which comprises a backbone (e.g., ResNet-18 [13]) and an MLP projector [4]. The latent representations of these images are denoted as $z_{i,\,c}=f(x_{i,\,c}),\,i=1,\,\cdots,\,n_{c}$. Then, an MLP predictor [11], denoted as $h$, transforms the $i$-th latent representation and matches it to the remaining latent representations. The prediction of $x_{i}$ for the remaining images is denoted as $p_{i,\,c}=h(f(x_{i,\,c}))$. The negative average cosine similarity of $x_{i,\,c}$ to other images is defined as:
$D(x_{i,\,c},\{x_{j,\,c}\}_{j\neq i}^{n_{c}})=-\frac{1}{n_{c}-1}\sum_{j\neq i}^{n_{c}}\frac{p_{i,\,c}}{\|p_{i,\,c}\|_{2}}\cdot\frac{z_{j,\,c}}{\|z_{j,\,c}\|_{2}},$ (1)
where $\|\cdot\|_{2}$ is $\ell_{2}$-norm. Finally, the overall loss is defined as:
$\mathcal{L}=\frac{1}{kn_{c}}\sum_{c=1}^{k}\sum_{i=1}^{n_{c}}D\left(x_{i,\,c},\,stopgrad\left(\{x_{j,\,c}\}_{j\neq i}^{n_{c}}\right)\right),$ (2)
where the stop-gradient operation ($stopgrad(\cdot)$) is an critical component to avoid model collapse when the contrastive model is trained exclusively with the positive pairs [5]. Specifically, $stopgrad(\{x_{j,\,c}\}_{j\neq i}^{n_{c}})$ entails treating each element in $\{z_{j,\,c}\}_{j\neq i}^{n_{c}}$ as a constant, ensuring no gradient flows from $z_{j,\,c}$ to $x_{j,\,c}$ in this term.
Actually, the proposed model is a more general form of simple Siamese (SimSiam) networks [5]. Consider a scenario where $k=n$ ($n$ is the number of images) and $n_{c}=2$, signifying that each image forms a cluster, with each cluster consists of the two augmented views of that image, the proposed model reverts to SimSiam. In this context, the proposed model leans towards grasping low-level semantics through the utilization of sample-level disparities among images, and the resulting representations are suboptimal for clustering tasks. In contrast, the latent representation learned by the generalized SimSiam (cluster-contrastive model), which takes into account the cluster-level distinctions among images, is more clustering-friendly.
Figure 3: CCFC architecture. The numbers indicate the order of the corresponding steps in each communication round.
3.2 Cluster-Contrastive Federated Clustering
Given a real-world dataset $X$ distributed among $m$ clients, i.e., $X=\bigcup_{l=1}^{m}X^{l}$. There are two straightforward methods for conducting contrastive learning in federated scenarios here: 1) Independent execution of contrastive learning by each client; 2) Integration of contrastive learning into the Federated Averaging (FedAvg) [21] framework. However, their performance is unsatisfactory. As shown in Fig. 1, one can observe: 1) The local model trained solely on local data exhibits poor performance. 2) Although the collaborative training can yield a better-performing global model, in the subsequent round of local updates, the limited information from the local data can lead to model regression. A similar problem of the model regression has also been observed in federated classification tasks, and it can be alleviated by introducing a model-contrastive loss to regularize the local model to refrain from deviating excessively from the global model during the local training [17].
Motivated by these, we propose a new federated clustering model, named as cluster-contrastive federated clustering (CCFC), which embeds the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term. As shown in Fig. 3 and Algorithm 1, CCFC involves three steps in each communication round: global information dissemination, local training and local information aggregation. Details are given below.
[…]
Local training. To acquire clustering-friendly representations using the cluster-contrastive model, each client $l$ first performs cluster assignment by labeling their local data with the index of the nearest global cluster centroid in the latent space, as defined in:
$\underset{c=\{1,\,\cdots,\,k\}}{\arg\min}\left\|f^{g}(x)-\eta^{g}_{c}\right\|_{2}.$ (3)
Then, each client $l$ trains the local model $w^{l}$ using their local data $X^{l}$ and the assignment results. To alleviate the model regression problem, we introduce a model-contrastive regularization term into Eq. 2. Finally, the revised loss function is denoted as:
$\mathcal{L_{R}}=\mathcal{L}+\frac{\lambda}{kn_{c}}\sum_{c=1}^{k}\sum_{i=1}^{n_{c}}R(w^{l}(x_{i,\,c}),\,w^{g}(x_{i,\,c})),$ (4)
where
$R(w^{l}(x_{i,\,c}),\,w^{g}(x_{i,\,c}))$ $=-\frac{p^{l}_{i,\,c}}{\|p^{l}_{i,\,c}\|_{2}}\cdot\frac{p^{g}_{i,\,c}}{\|p^{g}_{i,\,c}\|_{2}}$ (5)
$p^{l}_{i,\,c}$ $=h^{l}(f^{l}(x_{i,\,c})),$ (6)
$p^{g}_{i,\,c}$ $=h^{g}(f^{g}(x_{i,\,c})),$ (7)
and $\lambda$ is the tradeoff hyperparameter. By minimizing Eq. 4, the first term will encourage the local model to encode the local semantic structures discovered by the assignment result into the latent representation space, while the second one will encourage the local model $w^{l}$ to refrain from deviating excessively from the global model $w^{g}$. During the local training, the parameters of the global model remain fixed, while only those of the local model undergo updates.
After training, for the subsequent integration of local semantic information, each client first employs their respective local model to encode the local data into the latent representation space, and then use k-means (KM) to create $k$ local cluster centroids for storing the local semantic information.
Local information aggregation. First, each client uploads their trained local model and local cluster centroids to the server. Then, the server aggregates these local models into a new global model by computing their weighted average, as follows:
$w^{g}=\sum_{l=1}^{m}\frac{|X^{l}|}{|X|}w^{l}.$ (8)
And $k$ new global cluster centroids can be obtained by applying KM to these local cluster centroids.
So far, the whole communication loop of CCFC has been built up. And the final clustering result can be obtained by solving Eq. 3.
1 Initialize the global model $w^{g}$ and the global cluster centroids $\{\eta_{c}^{g}\}_{c=1}^{k}$.
[…]
3       Clients execute:
[…]
5             Global information dissemination:
6                 download $w^{g}$ and $\{\eta_{c}^{g}\}_{c=1}^{k}$ from server
7                 update the local model: $w^{l}=w^{g}$
8             Local training:
9                 group $X^{l}$ by solving Eq. 3
10                 train $w^{l}$ by minimizing Eq. 4
11                 mine $k$ local cluster centroids by KM
12             Local information aggregation:
13                 upload $w^{l}$ and the local cluster centroids
14                 to server
15
16       end for
17      Server executes:
18            Local information aggregation:
19               update $w^{g}$ by Eq. 8
20               update $\{\eta_{c}^{g}\}_{c=1}^{k}$ by applying KM to these
21               local cluster centroids.
22 end for
Cluster assignment: solve Eq. 3
Algorithm 1 CCFC
Table 1: NMI of clustering methods in different scenarios. For each comparison, the best result is highlighted in boldface.
Dataset $p$ Centralized setting Federated setting
KM FCM k-FED FFCM SDA-FC-KM SDA-FC-FCM PPFC-GAN CCFC(ours)
MNIST 0.0 0.5304 0.5187 0.5081 0.5157 0.5133 0.5141 0.6582 0.9236
0.25 0.4879 0.5264 0.5033 0.5063 0.6392 0.8152
0.5 0.4515 0.4693 0.5118 0.5055 0.6721 0.6718
0.75 0.4552 0.4855 0.5196 0.5143 0.7433 0.3611
1.0 0.4142 0.5372 0.5273 0.5140 0.8353 0.0766
Fashion-MNIST 0.0 0.6070 0.6026 0.5932 0.5786 0.5947 0.6027 0.6091 0.6237
0.25 0.5730 0.5995 0.6052 0.5664 0.5975 0.5709
0.5 0.6143 0.6173 0.6063 0.6022 0.5784 0.6023
0.75 0.5237 0.6139 0.6077 0.5791 0.6103 0.4856
1.0 0.5452 0.5855 0.6065 0.6026 0.6467 0.1211
CIFAR-10 0.0 0.0871 0.0823 0.0820 0.0812 0.0823 0.0819 0.1165 0.2449
0.25 0.0866 0.0832 0.0835 0.0818 0.1185 0.2094
0.5 0.0885 0.0870 0.0838 0.0810 0.1237 0.2085
0.75 0.0818 0.0842 0.0864 0.0808 0.1157 0.1189
1.0 0.0881 0.0832 0.0856 0.0858 0.1318 0.0639
STL-10 0.0 0.1532 0.1469 0.1468 0.1436 0.1470 0.1406 0.1318 0.2952
0.25 0.1472 0.1493 0.1511 0.1435 0.1501 0.1727
0.5 0.1495 0.1334 0.1498 0.1424 0.1432 0.2125
0.75 0.1455 0.1304 0.1441 0.1425 0.1590 0.1610
1.0 0.1403 0.1565 0.1477 0.1447 0.1629 0.0711
count - - - 0 2 1 0 6 11
Table 2: Kappa of clustering methods in different scenarios. For each comparison, the best result is highlighted in boldface.
Dataset $p$ Centralized setting Federated setting
KM FCM k-FED FFCM SDA-FC-KM SDA-FC-FCM PPFC-GAN CCFC(ours)
MNIST 0.0 0.4786 0.5024 0.5026 0.5060 0.4977 0.5109 0.6134 0.9619
0.25 0.4000 0.5105 0.4781 0.5027 0.5773 0.8307
0.5 0.3636 0.3972 0.4884 0.4967 0.6007 0.6534
0.75 0.3558 0.4543 0.4926 0.5021 0.6892 0.3307
1.0 0.3386 0.5103 0.5000 0.5060 0.7884 0.0911
Fashion-MNIST 0.0 0.4778 0.5212 0.4657 0.4974 0.4918 0.4918 0.4857 0.6411
0.25 0.5222 0.5180 0.4918 0.4918 0.4721 0.5261
0.5 0.4951 0.4974 0.4918 0.4918 0.4552 0.5929
0.75 0.4240 0.4995 0.4918 0.4918 0.4774 0.3945
1.0 0.3923 0.4672 0.4918 0.4918 0.5745 0.1434
CIFAR-10 0.0 0.1347 0.1437 0.1305 0.1439 0.1275 0.1283 0.1426 0.2854
0.25 0.1366 0.1491 0.1275 0.1376 0.1400 0.2281
0.5 0.1252 0.1316 0.1307 0.1411 0.1443 0.2214
0.75 0.1303 0.1197 0.1360 0.1464 0.1358 0.1214
1.0 0.1147 0.1237 0.1341 0.1494 0.1499 0.1047
STL-10 0.0 0.1550 0.1602 0.1390 0.1514 0.1533 0.1505 0.1557 0.1687
0.25 0.1361 0.1479 0.1448 0.1527 0.1611 0.1422
0.5 0.1505 0.1112 0.1377 0.1620 0.1415 0.1407
0.75 0.1256 0.1001 0.1513 0.1603 0.1813 0.1133
1.0 0.1328 0.1351 0.1527 0.1553 0.1868 0.0519
count - - - 0 1 0 2 7 10
4 Experiment
4.1 Experimental Setup
A universally adopted benchmark dataset for federated learning remains scarce at present, because the heterogeneity of the local data distributions among clients are unknown in real-world scenarios. Following [7, 31, 30], we simulate a range of federated scenarios by partitioning a real-world dataset into $k^{*}$ smaller subsets, each dedicated to a client, and adjusting the hyperparameter $p$ to scale the data heterogeneity for each client, where $k^{*}$ is the number of true clusters. Specifically, for the $l$-th client with $s$ images, $p\cdot s$ images are sampled from the $l$-th cluster, while the remaining $(1-p)\cdot s$ images are drawn randomly from the entire data. The hyperparameter $p$ varies from 0 (data is randomly distributed among $m$ clients) to 1 (each client forms a cluster). Our experiments are conducted on four real-world datasets: MNIST (70,000 images with 10 classes), Fashion-MNIST (70,000 images with 10 classes), CIFAR-10 (60,000 images with 10 classes), and STL-10 (13,000 images with 10 classes).
Recall that the proposed cluster-contrastive model requires the training of just two network modules: the encoder (a backbone plus an MLP projector) and an MLP predictor. For MNIST and Fashion-MNIST, the encoder includes three convolutional layers, while for CIFAR-10 and STL-10, we use ResNet-18 [13] as the backbone. For all the considered datasets, both the projector and predictor consist of two fully connected layers. The tradeoff hyperparameter $\lambda$ is set to 0.001, 1, 0.1 and 0.1 for MNIST, Fashion-MNIST, CIFAR-10 and STL-10, respectively. The latent representation dimensions are set to be 256, 64, 256 and 256, for the respective datasets.
The proposed method is trained using the Adam optimizer [16], implemented in the PyTorch [25], and the code will be made available.
4.2 Clustering Performance Comparison
We compare CCFC with five state-of-the-art methods, including k-FED [8], FFCM [27], SDA-FC-KM [31], SDA-FC-FCM [31] and PPFC-GAN [30]. To avoid excessive hyperparameter tuning, we use specific hyperparameter settings for each method across various simulated scenarios within the same dataset.
The clustering performance is assessed based on two metrics, NMI [28] and Kappa [18], as presented in Tab. 1 and Tab. 2, respectively. One can see that: 1) Although the two metrics exhibit different ranks in some federated scenarios, the advantages of the proposed method are quite significant in most scenarios. In some scenarios, its performance scores even double those of the best baseline methods. The most conspicuous case is observed on MNIST with the data heterogeneity $p=0$, where it achieves an NMI score surpassing PPFC-GAN by 0.2654 and a Kappa score exceeding 0.3485. 2) Although the synthetic data aided methods (SDA-FC-KM, SDA-FC-FCM and PPFC-GAN) are immune to, or even can benefit from, the data heterogeneity problem, the proposed method demonstrates a higher potential. 3) Representation learning is poised to become the primary catalyst for progress in federated clustering. The most related method to our proposed CCFC is k-fed, with the main distinction being the inclusion of a representation learning process in CCFC. Surprisingly, this learning process results in substantial NMI score improvements of up to 0.4155 and Kappa score enhancements of 0.4593 on the most conspicuous case. In the next subsection, we will present a more in-depth analysis to demonstrate the benefits of representation learning for federated clustering tasks.
(a) Raw pixels
(b) PPFC-GAN
(c) SCFC
(d) CCFC
Figure 4: t-SNE visualizations on MNIST with the data heterogeneity $p=0$ (best viewed in color and upon zooming in).
Table 3: Kappa of each client in one experiment. Imbalance ratio (IR) denotes the ration of the sample size of the largest class to the smallest class.
Client id IR Standalone models FedAvg models FedAvg models with model-contrastive learning
${}^{{\ddagger}}$ ${}^{{\ddagger}}$ ${}^{{\dagger}}$ ${}^{{\dagger}}$ SCFC CCFC
1 1.18 0.1286 0.4255 0.8243 0.8749 0.8970 0.9605
2 1.23 0.1774 0.2036 0.8297 0.8787 0.9046 0.9643
3 1.39 0.1685 0.0517 0.8273 0.8783 0.9027 0.9611
4 1.21 0.2751 0.1847 0.8239 0.8714 0.8987 0.9617
5 1.24 0.2134 0.0926 0.8179 0.8689 0.8976 0.9609
6 1.23 0.1862 0.1892 0.8201 0.8696 0.8987 0.9625
7 1.19 0.1876 0.1681 0.8241 0.8722 0.8995 0.9624
8 1.19 0.3407 0.3741 0.8330 0.8750 0.9030 0.9611
9 1.18 0.2810 0.2936 0.8190 0.8687 0.8983 0.9616
10 1.53 0.1940 0.2727 0.8214 0.8686 0.9000 0.9630
Table 4: MNIST test accuracy (%) for kNN on different representation spaces.
k 3 5 7 9 100 1000 7000
Raw pixels 97.05 96.88 96.94 96.59 94.40 87.31 70.16
PPFC-GAN 95.26 95.20 95.25 95.25 92.73 85.37 71.19
SCFC 98.11 98.23 98.25 98.21 97.54 96.30 94.09
CCFC 98.25 98.33 98.26 98.32 97.86 97.09 97.09
4.3 Representation Learning
We claim that the advantages of PPFC-GAN and CCFC tasks stem from the fact that the learned representations are more clustering-friendly. To substantiate this, we qualitatively and quantitatively investigate the latent representations learned by them on MNIST with the data heterogeneity $p=0$. Moreover, to demonstrate the efficacy of the cluster-contrastive learning, we construct an additional baseline method, named sample-contrastive federated clustering (SCFC), achieved by reverting the cluster-contrastive model to SimSiam. This signifies that, throughout the local training process, SCFC abstains from leveraging the semantic structures discovered by clustering.
Qualitative analysis. As shown in Fig. 4, one can observer that: 1) In the original data space, there are 10 discernible modes, yet many samples from different categories are mixed up together, necessitating representation learning. 2) Representation learning through contrastive learning significantly alleviates the mixing problem, and the cluster-contrastive learning yields more compact representations. 3) Interestingly, the latent representations learned by PPFC-GAN seem to exacerbate the mixing problem, implying that the advantages of PPFC-GAN may not be attributed to the acquisition of more clustering-friendly representations. Hence, to thoroughly figure out this matter, further quantitative analysis is necessary.
Quantitative analysis. To offer more compelling evidence that representation learning can mitigate the mixing problem in the original data space, we run the k-nearest neighbors (KNN) classifier in different representation spaces. The better the test performance, the fewer samples from different categories among the $k$ neighbors, suggesting a reduced severity of the mixing problem.
As shown in Tab. 4, one can find that: 1) The test performance in the latent spaces is comparable to, or in some cases, inferior to that in the original data space for small values of $k$. However, as $k$ values increase, the merits of latent representation learning progressively come to fruition by mitigating the mixing problem. 2) The proposed method demonstrates superior effectiveness and robustness, better preserving the inherent semantic structure of the data.
(a) NMI
(b) Kappa
Figure 5: Sensitivity of CCFC to $\lambda$ on MNIST with different data heterogeneity.
Figure 6: Sensitivity of CCFC to the device failures on MNIST with different data heterogeneity.
4.4 Ablation Study
CCFC comprises three key components: cluster-contrastive learning, FedAvg and the model-contrastive regularization term. To validate the effectiveness of each component, we devise four additional baselines within two distinct federated clustering frameworks:
1) Standalone models. In this set of baselines, each client independently performs sample-contrastive learning (denoted as ${}^{{\ddagger}}$) or cluster-contrastive learning (denoted as ${}^{{\ddagger}}$).
2) FedAvg models. In this set of baselines, each client interactively performs sample-contrastive learning (denoted as ${}^{{\dagger}}$) or cluster-contrastive learning (denoted as ${}^{{\dagger}}$) within the FedAvg framework, without employing the model-contrastive regularization term.
Tab. 3 shows that: 1) The complete model exhibits the most favorable clustering performance, implying that the removal of any component can result in performance degradation. 2) The most substantial improvements arise from the collaborative training among clients, while cluster-contrastive learning and the model-contrastive regularization term can yield further performance improvements.
4.5 Hyperparameter Sensitivity Analysis
Although we have previously validated the efficacy of the model-contrastive regularization term, the relative importance of this term compared to the cluster-contrastive loss remains unexplored. To this end, we conduct a series of CCFC experiments with different $\lambda$ on MNIST.
Fig. 5 reveals that: 1) The clustering performance remains robust across a wide range of $\lambda$ under a fixed federated heterogeneous scenario. 2) The optimal value of $\lambda$ differs across different federated scenarios. Consequently, the performance of CCFC within Tab. 1 and Tab. 2 is somewhat underestimated, as we used the same $\lambda$ for different federated scenarios of the same dataset to prevent excessive hpyerparameter tuning. 3) The sensitivity of the model to data heterogeneity is more pronounced compared to $\lambda$.
4.6 Device Failures
In real-world scenarios, client devices may occasionally fail to connect with the server during the training process due to wireless network fluctuations, energy constraints, etc. As a result, specific data characteristics from the failed devices may be lost, resulting in poor and unrobust performance. Hence, it is essential to investigate the sensitivity of CCFC to device failures from a practical viewpoint.
Following [31, 30], we simulate different disconnection scenarios by scaling the disconnection rate, which measures the proportion of disconnected clients among all clients. Throughout the training process, only the connected clients are involved. As shown in Fig. 6, while CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
In summary: 1) The representations acquired through the proposed cluster-contrastive learning model are more clustering-friendly. 2) Benefiting from the more clustering-friendly representations, CCFC demonstrates superior performance. 3) The considered key components within CCFC are all essential for its superiority. 4) The clustering performance remains robust across a wide range of the tradeoff hyperparameter $\lambda$. 5) While CCFC exhibits greater sensitivity to device failures compared to the baselines, it remains the best performance in most cases.
5 Conclusion
In this work, we first illustrate the necessity of representation learning in federated clustering tasks through a motivational example, and then highlight the limitations of straightforward extensions from centralized models. To handle these, we propose a cluster-contrastive model for learning more clustering-friendly representations, and introduce a simple yet effective federated clustering method by embedding the cluster-contrastive model into the FedAvg framework with the model-contrastive regularization term.
Comprehensive experiments demonstrate the superiority of the cluster-contrastive model in learning clustering-friendly representations, as well as the excellence of the proposed clustering method in terms of clustering performance and handling device failures. We hope our work will draw the community’s focus to the fundamental role of representation learning in federated clustering, and lay a foundation for future advancements within the community.